# No-prior Bayes reIMagined: probabilistic approximations of possibilistic inferential models

### v2

@@ B.1 Theory
The last line in the above display can be recognized as the probability-to-possibility transform applied to the Bayesian/fiducial posterior distribution. This identifies the posterior distribution $\mathsf{Q}_{x}^{\text{\sc b}}$ as the inner probabilistic approximation of the possibilistic IM $\overline{\Pi}_{x}=\overline{\Pi}_{g|u}$, and hence a new/different proof of the main result in Martin, 2023a , which is also Theorem 3 above.
B.2 Directional data example
@@
Here $G$ is the angle the vector $(\bar{C},\bar{S})$ makes with the horizontal axis and $U$ is the Euclidean length of $(\bar{C},\bar{S})$. It’s clear that $(G,U)$ is a bijection of the minimal sufficient statistic $Y$, so write $t(X)=(G,U)$. I’m using the notation $G$ and $U$ to align with that introduced in the previous section. That is, there is a group of transformations $\mathscr{G}$ that corresponds to rotations of points in the plane; technically, $\mathscr{G}$ is the special orthogonal group $\text{SO}(2)$ consisting of orthogonal matrices with unit determinant. Then $G$ can be interpreted as both an angle or as a rotation by that angle. Similarly, the interior of the unit circle can be partitioned into distinct concentric circles, which correspond to orbits, and $U$ determines which of these orbits $Y$ sits on.
The details here align with the general setup above, so I can proceed to carry out the fiducial (or default-prior Bayes) and IM analyses. It’s a standard result (e.g., Mardia and Jupp, 2000, Eq. 4.5.5) that the joint density for $t(X)=(G,U)$ factors as
@@
=$=\{2\pi I_{0}(\kappa u)\}^{-1}\exp\{\kappa u\cos(g-\theta)\}\,p(u),$
@@
Appendix C Details from Section 4.2
C.1 Setup and regularity conditions
@@
The condition in (21) is often described as $\theta\mapsto s_{\theta}$ being differentiable in quadratic mean. Note that this condition does not require the square-root density to actually be differentiable at $\theta$, only that it be locally “linearizable” in an average sense, like a differentiable-at-$\theta$ function would be. The classical Cramér conditions assume more than two continuous derivatives, so (21), which does not even require existence of a first derivative, is significantly weaker; sufficient conditions for (21) are given in van der Vaart, (1998, Lemma 7.6). Then the score function ˙$\dot{\ell}_{\theta}(x)$ is defined in terms of $\dot{s}_{\theta}(x)$ as
[…]
and it can be shown that ˙$\int\dot{\ell}_{\theta}(x)\,\mathsf{P}_{\theta}(dx)=0$ for each $\theta$. Moreover, Condition A3 above implies non-singularity of the Fisher information matrix $I_{\theta}=\int\dot{\ell}_{\theta}(x)\ell_{\theta}(x)^{\top}\,\mathsf{P}_{\theta}(dx)$ for each $\theta\in\mathbb{T}$. Bickel et al., (1998, Prop. 2.1.1) provide sufficient conditions for A1–A3.
@@ C.2 Proof of Theorem 5
[…]
@@
$\pi_{X^{n}}(\theta)\in(\alpha_{1},\alpha_{2})\quad\text{and}\quad\gamma_{X^{n}}(\theta)>\alpha_{2}\quad\text{for all $\theta$ in that subset;}$;
of course, the last inequality could be replaced by “$\gamma_{X^{n}}(\theta)<\alpha_{1}$ and the positions of $\pi_{X^{n}}$ and $\gamma_{X^{n}}$ in the above display could be reversed, but these changes don’t affect the argument that follows. What matters is that there exists a point $\theta$—depending implicitly on data $X^{n}$, on $H$, and on $\varepsilon$—such that
@@
What’s not fully satisfactory about the above conclusion is that the positive bound “$\alpha_{2}-\alpha_{1}$” is only implicit. An important point that the above argument reveals is that any differences in how $\mathsf{Q}_{X^{n}}^{\star}$ and $\mathsf{G}_{X^{n}}$ assign probability will manifest through the contour level sets. Since these are also the sets that determine the distributions themselves, it’s without loss of generality to consider hypotheses $H$ that take the form of lower level sets of either of the defining contours. For concreteness, consider
@@
Putting everything together, $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$ implies existence of a set $H$ for which $|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon$. This implies a non-negligible difference in the contours so, without loss of generality, the problematic $H$—whose existence is guaranteed—can be taken as a lower level set corresponding to one of the contours. When $H$ is such a lower level set, $|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon$ implies that $|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|>\varepsilon$ for all $\theta$ on the boundary of $H$. Since the latter event has vanishing $\mathsf{P}_{\Theta}$-probability by Theorem 4, so must the event $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$, thus proving Theorem 5.
Appendix D Details from Section 4.3
@@
Appendix E Computational details from Section 5
Here I review some of the relevant details presented in Cella and Martin, (2024) on the choice of $\xi=\xi(x,\alpha)$ in Section 5. In this review I’ll focus on use of a Gaussian variational family, which is motivated by the asymptotic results in Section 4.2, but similar developments can surely be made with other distributional families. The key features of the Gaussian are that its credible sets can be described in closed-form (ellipsoids) and that it’s straightforward to sample on the boundary of these credible sets.
Let $\mathsf{G}_{x}^{\xi}$ denote the Gaussian distribution ${\sf N}_{D}(\hat{\theta}_{x},(J_{x}^{\xi})^{-1})$ with mean equal to the maximum likelihood estimator and covariance matrix the inverse of the $\xi$-adjusted observed Fisher information matrix in (15). Recall that $\xi$ scales the eigenvalues of $J_{x}^{\xi}$, in particular, the eigenvalues are decreasing elementwise in $\xi$. The corresponding Gaussian credible/confidence sets are
$C_{\alpha}^{\xi}(x)=\{\theta:(\theta-\hat{\theta}_{x})^{\top}J_{x}^{\xi}(\theta-\hat{\theta}_{x})\leq\chi_{D,1-\alpha}^{2}\},\quad\alpha\in[0,1].$
The intuition is that $C_{\alpha}^{\xi}(x)$ can be appropriately inflated or deflated by varying $\xi$, and this is relevant because the Gaussian approximation $\mathsf{G}_{x}^{\xi}$ assigns probability at least $1-\alpha$ to the original IM’s $\alpha$-cut $C_{\alpha}(x)$ if $C_{\alpha}^{\xi}(x)\supseteq C_{\alpha}(x)$ or, equivalently, if
$\sup_{\theta\not\in C_{\alpha}^{\xi}(x)}\pi_{x}(\theta)\leq\alpha.$
Since the contour $\pi_{x}$ is itself approximately Gaussian (Martin and Williams, 2025) and the maximum likelihood estimator $\hat{\theta}_{x}$, also the mode of $\pi_{x}$, is in $C_{\alpha}^{\xi}(x)$, the above supremum should be attained on the boundary $\partial C_{\alpha}^{\xi}(x)$. Moreover, since equality in the above display implies a near-perfect match between the IM’s and the posited Gaussian $\alpha$-cuts, a reasonable goal is to find a root to the function
$g_{\alpha}(\xi):=\max_{\theta\in\partial C_{\alpha}^{\xi}(x)}\pi_{x}(\theta)-\alpha.$
Design of an iterative algorithm to find this root requires care, primarily because evaluating $\pi_{x}$ is expensive; so the goal is to evaluate $g_{\alpha}(\xi)$ with as few $\pi_{x}$ evaluations as possible. Cella and Martin, (2024) propose to represent the boundary of $C_{\alpha}^{\xi}(x)$ by $2D$-many vectors
$\vartheta_{s}^{\xi,\pm}:=\hat{\theta}_{x}\pm\bigl{\{}\chi_{D,1-\alpha}^{2}\,\xi_{s}\,/\,\lambda_{s}\}^{1/2}\,e_{s},\quad s=1,\ldots,d,$ (23)
@@ Appendix E Computational details from Section 5
$\hat{g}_{\alpha,s}(\xi)=\max\{\pi_{x}(\vartheta_{s}^{\xi,+}),\pi_{x}(\vartheta_{s}^{\xi,-})\}-\alpha,\quad s=1,\ldots,D.$ (24)
At least intuitively, negative and positive $\hat{g}_{\alpha,s}(\xi)$ indicate that the $\alpha$-cut $C_{\alpha}^{\xi}(x)$ is too large and too small, respectively in the $e_{s}$-direction. From here, Cella and Martin, (2024) suggest applying a stochastic approximation algorithm—see Robbins and Monro, (1951) and Kushner and Yin, (2003)—to construct a sequence $(\xi^{(t)}:t\geq 1)$ of $D$-vectors that converges to a root of $\hat{g}_{\alpha}$ and, hence, an approximate root of $g_{\alpha}$. For an initial guess $\xi^{(0)}$, the specific sequence is defined as
$\xi_{s}^{(t+1)}=\xi_{s}^{(t)}+w_{t+1}\,\hat{g}_{\alpha,s}(\xi^{(t)}),\quad s=1,\ldots,D,\quad t\geq 0,$
@@
These steps are iterated until (practical) convergence is achieved, and the limit is what I called $\xi(x,\alpha)$ in Section 5. The basic steps are outlined in Algorithm 1, but I refer to Cella and Martin, (2024) for further details and discussion.
@@
initialize: $\alpha$-level, guess $\xi^{(0)}$, step size sequence $(w_{t})$, and threshold $\varepsilon>0$;
@@
       construct the representative points $\{\vartheta_{s}^{\xi^{(t)},\pm}:s=1,\ldots,D\}$ as in (23);
       evaluate $\hat{g}_{\alpha,s}(\xi_{s}^{(t)})$ for $s=1,\ldots,D$ as in (24);
       update $\xi_{s}^{(t+1)}=\xi_{s}^{(t)}\pm w_{t+1}\,\hat{g}_{\alpha,s}(\xi_{s}^{(t)})$ for $s=1,\ldots,D$;
       if $\max_{s}|\xi_{s}^{(t+1)}-\xi_{s}^{(t)}|<\varepsilon$ then
             $\xi(z,\alpha)=\xi^{(t+1)}$;
@@
return $\xi(x,\alpha)$;
Algorithm 1 Determining $\xi(x,\alpha)$—from Cella and Martin, (2024)
Appendix F Additional examples
F.1 Bivariate normal correlation
A challenging one-parameter inference problem is that where $X=(X_{1},\ldots,X_{n})$ are iid, with $X_{i}=(X_{1i},X_{2i})$ a bivariate normal random vector with zero means, unit standard deviations, and correlation $\Theta\in\mathbb{T}=(-1,1)$ to be inferred. Ironically, the known mean and variance case is more difficult—it’s a curved exponential family so the minimal sufficient statistic is not complete and there are various ancillary statistics available to condition on (e.g., Basu, 1964). How this curvature affects asymptotic inference is detailed in Reid, (2003). A likelihood-based possibilistic IM, which is exactly valid for all sample sizes and asymptotically efficient, is easy to construct in theory (Martin, 2024a , Example 2), but naive computation is relatively expensive because there’s no pivot structure and no closed-form expression for the maximum likelihood estimator. This motivates the faster, approximate solution suggested in Martin, (2025) and Section 5 above, based on the variational approximation strategy developed in Cella and Martin, (2024).
A relevant feature of the exact IM solution is that, when $\Theta$ is relatively close the extremes, i.e., $\Theta\approx\pm 1$, the IM’s possibility contour tends to be very asymmetric. Consequently, approximation suggested in Martin, (2025) and Section 5 above, which are based on stitching together symmetric, Gaussian credible intervals, tend to be a bit conservative. Other asymmetric distributions, e.g., skewed normal (e.g., Azzalini, 1985), could potentially replace the Gaussian recommendation made in Martin, (2025), but that’s beyond the scope of the present paper. I’m mentioning this point here because the same conservative behavior will be observed below with the inner probabilistic approximation; this isn’t a shortcoming of the inner probabilistic approximation itself, but of the choice to bound the IM’s asymmetric level sets using symmetric Gaussian credible limits.

### v3

@@ C.1 Theory
The last line in the above display can be recognized as the probability-to-possibility transform applied to the Bayesian/fiducial posterior distribution. This identifies the posterior distribution $\mathsf{Q}_{x}^{\rho}$ as an inner probabilistic approximation of the possibilistic IM $\mathsf{\Pi}_{x}=\mathsf{\Pi}_{g|u}$. This provides a new/different proof of the main result in Martin, 2023a , of which Theorem 3 above is a generalization.
As mentioned above, there are cases in which the group $\mathscr{G}$ with respect to which the model is invariant is not unique. A key example of this is in the case of Gaussian models, which are invariant with respect to linear (or affine) transformations $gx=Ax+b$, indexed by pairs $(A,b)$ consisting of a matrix $A$ and a vector $b$. Since the group indexed by all such matrices $A$ is “too large” (in the sense that the underlying group is not amenable), it’s natural to restrict the transformations to those indexed by certain sub-classes of matrices, corresponding to one or another lower-dimensional matrix factorization. It turns out that the right Haar prior depends on which class of matrix factorizations is used. For example, in the bivariate normal model indexed by means $\mu_{1}$ and $\mu_{2}$, standard deviations $\sigma_{1}$ and $\sigma_{2}$, and correlation $\rho$, there are two right Haar priors (Sun and Berger, 2007, Sec. 5), with respective density functions
$(\mu_{1},\mu_{2},\sigma_{1},\sigma_{2},\rho)\mapsto\frac{1}{\sigma_{2}^{2}(1-\rho^{2})}\quad\text{and}\quad(\mu_{1},\mu_{2},\sigma_{1},\sigma_{2},\rho)\mapsto\frac{1}{\sigma_{1}^{2}(1-\rho^{2})}.$
Of course, these two distinct priors lead to (slightly) different Bayes solutions with different statistical properties. Some might find this troubling: since the model is what determines the statistical inference problem, not the choice of group representation, it’s problematic if the Bayesian solution depends on a somewhat ad hoc choice between one group/right Haar prior or another.
In the construction of the IM solution described above, the dependence on the group structure is implicit: it’s in the identification of the $(G,U)$ components, where $G$ is an equivariant function of $X$ and $U$ is a (maximal) invariant function of $X$ which, of course, depend on $\mathscr{G}$. So, in general, different groups with respect to which the model is invariant would suggest a different “(G,U)(G,U)” representation. Since this representation determines how the conditioning is carried out in (22), this IM solution also depends on the particular choice of group $\mathscr{G}$. But I want to emphasize that conditioning on an ancillary statistic isn’t required for the IM construction; this is only recommended when an appropriate ancillary statistic is available. If the group isn’t unique, then there are multiple ancillary statistics available, yielding distinct conditional IMs. There are some natural ways one might remove the group-specific conditional IM’s undesirable dependence on the choice of group, but this is not really relevant to the present paper, so I’ll save this investigation for future work; see, also, the next paragraph. Alternatively, when there are multiple ancillary statistics, arguably none of them are fully “appropriate.” In that case, one might prefer to ignore conditioning and proceed with the basic likelihood-based IM construction as in Equation (2) of the main text. A corresponding inner probabilistic approximation retains the desirable probability matching and Bernstein–von Mises properties, but whether it corresponds to any Bayesian posterior distribution is an interesting open question.
To conclude this discussion about non-unique groups, let me consider the aforementioned Gaussian case specifically. There, the maximum likelihood estimator is a complete and sufficient statistic, so, by Basu’s theorem, whether one conditions on an ancillary statistic or not has no effect on the IM solution. Consequently, the same argument used above to prove Theorem 3 applied to the present case gives
$\pi_{x}(\theta)=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\}=\mathsf{Q}_{x}^{\rho}\{q_{x}^{\rho}(\Theta)\leq q_{x}^{\rho}(\theta)\},\quad\theta\in\mathbb{T},$
for any right Haar prior $\rho$. Therefore, in the multivariate Gaussian case, it follows that all right Haar prior Bayes solutions are inner probabilistic approximations of the IM with contour as defined in Equation (2) of the main text. Compare this with the results in Eaton and Sudderth, (2012) which imply, for example, that highest posterior density regions corresponding to different right Haar priors in these multivariate normal models are exactly probability matching. Outside the Gaussian family, completeness typically doesn’t hold, so I currently don’t know if the above claim concerning “all right Haar prior Bayes solutions” applies more broadly.
C.2 Directional data example
@@
Here $G$ is the angle the vector $(\bar{C},\bar{S})$ makes with the horizontal axis and $U$ is the Euclidean length of $(\bar{C},\bar{S})$. It’s clear that $(G,U)$ is a bijection of the minimal sufficient statistic $Y$, so write $t(X)=(G,U)$. I’m using the notation $G$ and $U$ to align with that introduced in the previous subsection. That is, there’s a group of transformations $\mathscr{G}$ that corresponds to rotations of points in the plane; technically, $\mathscr{G}$ is the special orthogonal group $\text{SO}(2)$ consisting of orthogonal matrices with unit determinant. Then $G$ can be interpreted as both an angle or as a rotation by that angle. Similarly, the interior of the unit circle can be partitioned into distinct concentric circles, which correspond to orbits, and $U$ determines which of these orbits $Y$ sits on.
The details here align with the general setup above, so I can proceed to carry out the fiducial or default-prior Bayes and IM analyses. It’s a standard result (e.g., Mardia and Jupp, 2000, Eq. 4.5.5) that the joint density for $t(X)=(G,U)$ factors as
@@
$=\{2\pi I_{0}(\kappa u)\}^{-1}\exp\{\kappa u\cos(g-\theta)\}\,p(u),$
@@
Appendix D Details from Section 4.2
D.1 Setup and regularity conditions
@@
The condition in (23) is often described as $\theta\mapsto s_{\theta}$ being differentiable in quadratic mean. Note that this condition does not require the square-root density to actually be differentiable at $\theta$, only that it be locally “linearizable” in an average sense, like a differentiable-at-$\theta$ function would be. The classical Cramér conditions assume more than two continuous derivatives, so (23), which does not even require existence of a first derivative, is significantly weaker; sufficient conditions for (23) are given in van der Vaart, (1998, Lemma 7.6). Then the score function $\dot{\ell}_{\theta}(x)$ is defined in terms of $\dot{s}_{\theta}(x)$ as
$\dot{\ell}_{\theta}(x)=\frac{2\dot{s}_{\theta}(x)}{s_{\theta}(x)}\,1\{s_{\theta}(x)>0\},$
and it can be shown that $\int\dot{\ell}_{\theta}(x)\,\mathsf{P}_{\theta}(dx)=0$ for each $\theta$. Moreover, Condition A3 above implies non-singularity of the Fisher information matrix $I_{\theta}=\int\dot{\ell}_{\theta}(x)\ell_{\theta}(x)^{\top}\,\mathsf{P}_{\theta}(dx)$ for each $\theta\in\mathbb{T}$. Bickel et al., (1998, Prop. 2.1.1) provide sufficient conditions for A1–A3.
@@ D.2 Proof of Theorem 5
$\mathsf{Q}_{X^{n}}^{\star}(\cdot)=\int_{0}^{1}\mathsf{K}_{X^{n},\pi}^{\alpha}(\cdot)\,d\alpha\quad\text{and}\quad\mathsf{G}_{X^{n}}(\cdot)=\int_{0}^{1}\mathsf{K}_{X^{n},\gamma}^{\alpha}(\cdot)\,d\alpha.$
@@
$\pi_{X^{n}}(\theta)\in(\alpha_{1},\alpha_{2})\quad\text{and}\quad\gamma_{X^{n}}(\theta)>\alpha_{2}\quad\text{for all $\theta$ in that subset;}$
of course, the last inequality could be replaced by $\gamma_{X^{n}}(\theta)<\alpha_{1}$ and the positions of $\pi_{X^{n}}$ and $\gamma_{X^{n}}$ in the above display could be reversed, but these changes don’t affect the argument that follows. What matters is that there exists a point $\theta$—depending implicitly on data $X^{n}$, on $H$, and on $\varepsilon$—such that
@@
What’s not fully satisfactory about the above conclusion is that the positive bound $\alpha_{2}-\alpha_{1}$” is only implicit. An important point that the above argument reveals is that any differences in how $\mathsf{Q}_{X^{n}}^{\star}$ and $\mathsf{G}_{X^{n}}$ assign probability will manifest through the contour level sets. Since these are also the sets that determine the distributions themselves, it’s without loss of generality to consider hypotheses $H$ that take the form of lower level sets of either of the defining contours. For concreteness, consider
@@
Putting everything together, $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$ implies existence of a set $H$ for which $|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon$. This implies a non-negligible difference in the contours so, without loss of generality, the problematic HH—whose existence is guaranteed—can be taken as a lower level set corresponding to one of the contours. When $H$ is such a lower level set, $|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)|>\varepsilon$ implies that $|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)|>\varepsilon$ for all $\theta$ on the boundary of $H$. Since the latter event has vanishing $\mathsf{P}_{\Theta}$-probability by Theorem 4, so must the event $d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}})>\varepsilon$, thus proving Theorem 5.
Appendix E Details from Section 4.3
@@
Appendix F Computational details from Section 5
Here I review some of the relevant details presented in Cella and Martin, (2025) on the choice of $\sigma=\sigma(x,\alpha)$ in Section 5. In this review I’ll focus on use of a Gaussian variational family, which is motivated by the asymptotic results in Section 4.2, but similar developments can surely be made with other distributional families. The key features of the Gaussian are that its credible sets can be described in closed-form (ellipsoids) and that it’s straightforward to sample on the boundary of these credible sets.
Let $\mathsf{G}_{x}^{\sigma}$ denote the Gaussian distribution ${\sf N}_{D}(\hat{\theta}_{x},(J_{x}^{\sigma})^{-1})$ with mean equal to the maximum likelihood estimator and covariance matrix the inverse of the $\sigma$-adjusted observed Fisher information matrix in (15). Recall that $\sigma$ scales the eigenvalues of $J_{x}^{\sigma}$, in particular, the eigenvalues are decreasing elementwise in $\sigma$. The corresponding Gaussian credible/confidence sets are
$C_{\alpha}^{\sigma}(x)=\{\theta:(\theta-\hat{\theta}_{x})^{\top}J_{x}^{\sigma}(\theta-\hat{\theta}_{x})\leq\chi_{D,1-\alpha}^{2}\},\quad\alpha\in[0,1].$
The intuition is that $C_{\alpha}^{\sigma}(x)$ can be appropriately inflated or deflated by varying $\sigma$, and this is relevant because the Gaussian approximation $\mathsf{G}_{x}^{\sigma}$ assigns probability at least $1-\alpha$ to the original IM’s $\alpha$-cut $C_{\alpha}(x)$ if $C_{\alpha}^{\sigma}(x)\supseteq C_{\alpha}(x)$ or, equivalently, if
$\sup_{\theta\not\in C_{\alpha}^{\sigma}(x)}\pi_{x}(\theta)\leq\alpha.$
Since the contour $\pi_{x}$ is itself approximately Gaussian (Martin and Williams, 2025) and the maximum likelihood estimator $\hat{\theta}_{x}$, also the mode of $\pi_{x}$, is in $C_{\alpha}^{\sigma}(x)$, the above supremum should be attained on the boundary $\partial C_{\alpha}^{\sigma}(x)$. Moreover, since equality in the above display implies a near-perfect match between the IM’s and the posited Gaussian $\alpha$-cuts, a reasonable goal is to find a root to the function
$g_{\alpha}(\sigma):=\max_{\theta\in\partial C_{\alpha}^{\sigma}(x)}\pi_{x}(\theta)-\alpha.$
Design of an iterative algorithm to find this root requires care, primarily because evaluating $\pi_{x}$ is expensive; so the goal is to evaluate $g_{\alpha}(\sigma)$ with as few $\pi_{x}$ evaluations as possible. Cella and Martin, (2025) propose to represent the boundary of $C_{\alpha}^{\sigma}(x)$ by $2D$-many vectors
$\vartheta_{s}^{\sigma,\pm}:=\hat{\theta}_{x}\pm\bigl\{\chi_{D,1-\alpha}^{2}\,\sigma_{s}\,/\,\lambda_{s}\}^{1/2}\,e_{s},\quad s=1,\ldots,d,$ (25)
@@ Appendix F Computational details from Section 5
$\hat{g}_{\alpha,s}(\sigma)=\max\{\pi_{x}(\vartheta_{s}^{\sigma,+}),\pi_{x}(\vartheta_{s}^{\sigma,-})\}-\alpha,\quad s=1,\ldots,D.$ (26)
At least intuitively, negative and positive $\hat{g}_{\alpha,s}(\sigma)$ indicate that the $\alpha$-cut $C_{\alpha}^{\sigma}(x)$ is too large and too small, respectively in the $e_{s}$-direction. From here, Cella and Martin, (2025) suggest applying a stochastic approximation algorithm—see Robbins and Monro, (1951) and Kushner and Yin, (2003)—to construct a sequence $(\sigma^{(t)}:t\geq 1)$ of $D$-vectors that converges to a root of $\hat{g}_{\alpha}$ and, hence, an approximate root of $g_{\alpha}$. For an initial guess $\sigma^{(0)}$, the specific sequence is defined as
$\sigma_{s}^{(t+1)}=\sigma_{s}^{(t)}+w_{t+1}\,\hat{g}_{\alpha,s}(\sigma^{(t)}),\quad s=1,\ldots,D,\quad t\geq 0,$
@@
These steps are iterated until (practical) convergence is achieved, and the limit is what I called $\sigma(x,\alpha)$ in Section 5. The basic steps are outlined in Algorithm 1, but I refer to Cella and Martin, (2025) for further details and discussion.
@@
initialize: $\alpha$-level, guess $\sigma^{(0)}$, step size sequence $(w_{t})$, and threshold $\varepsilon>0$;
@@
   construct the representative points $\{\vartheta_{s}^{\sigma^{(t)},\pm}:s=1,\ldots,D\}$ as in (25);
   evaluate $\hat{g}_{\alpha,s}(\sigma_{s}^{(t)})$ for $s=1,\ldots,D$ as in (26);
   update $\sigma_{s}^{(t+1)}=\sigma_{s}^{(t)}\pm w_{t+1}\,\hat{g}_{\alpha,s}(\sigma_{s}^{(t)})$ for $s=1,\ldots,D$;
 if $\max_{s}|\sigma_{s}^{(t+1)}-\sigma_{s}^{(t)}|<\varepsilon$ then
    $\sigma(z,\alpha)=\sigma^{(t+1)}$;
@@
return $\sigma(x,\alpha)$;
Algorithm 1 Determining $\sigma(x,\alpha)$—from Cella and Martin, (2025)
Appendix G Additional examples
G.1 Bivariate normal correlation
A challenging one-parameter inference problem is that where $X=(X_{1},\ldots,X_{n})$ are iid, with $X_{i}=(X_{1i},X_{2i})$ a bivariate normal random vector with zero means, unit standard deviations, and correlation $\Theta\in\mathbb{T}=(-1,1)$ to be inferred. Ironically, the known mean and variance case is more difficult—it’s a curved exponential family so the minimal sufficient statistic is not complete and there are various ancillary statistics available to condition on (e.g., Basu, 1964). How this curvature affects asymptotic inference is detailed in Reid, (2003). A likelihood-based possibilistic IM, which is exactly valid for all sample sizes and asymptotically efficient, is easy to construct in theory (Martin, 2024a , Example 2), but naive computation is relatively expensive because there’s no pivot structure and no closed-form expression for the maximum likelihood estimator. This motivates the faster, approximate solution suggested in Martin, 2025a and Section 5 above, based on the variational approximation strategy developed in Cella and Martin, (2025).
A relevant feature of the exact IM solution is that, when $\Theta$ is relatively close the extremes, i.e., $\Theta\approx\pm 1$, the IM’s possibility contour tends to be very asymmetric. Consequently, approximation suggested in Martin, 2025a and Section 5 above, which are based on stitching together symmetric, Gaussian credible intervals, tend to be a bit conservative. Other asymmetric distributions, e.g., skewed normal (e.g., Azzalini, 1985), could potentially replace the Gaussian recommendation made in Martin, 2025a , but that’s beyond the scope of the present paper. I’m mentioning this point here because the same conservative behavior will be observed below with the inner probabilistic approximation; this isn’t a shortcoming of the inner probabilistic approximation itself, but of the choice to bound the IM’s asymmetric level sets using symmetric Gaussian credible limits.